OpenStudy (anonymous):

I am new to the group and I am studying trig identities. I have a problem that asks me to express the sin function as an expression of cot. Can someone help?

OpenStudy (anonymous):

More information: sin\[\sin \theta = \cos \theta \div \cot \theta\]

OpenStudy (anonymous):

wt do we hve to do in this prob?

OpenStudy (anonymous):

The problem is express Sin as a function of Cot

OpenStudy (anonymous):

that would be a miracle

OpenStudy (anonymous):

i suppose you could write \[\cot(x)=\pm\frac{\sqrt{1-\sin^2(x)}}{\sin(x)}\]

OpenStudy (anonymous):

oh worry you wanted sin in terms of cotangent and i gave it the other way

OpenStudy (anonymous):

Thanks for noting that! It was going to be my next question.

OpenStudy (anonymous):

der wuld b couple f more ways to write it.....wt say satellite73?

OpenStudy (anonymous):

good. glad we got one done at least.

OpenStudy (anonymous):

strikes me as silly. but what do i know?

OpenStudy (anonymous):

Trig identities are a little strange I agree, but it is still one of the key exercises. I guess I am still wondering about the steps in between. Thanks for trying.